1. The daily rainfall in Dublin (measured in millimeters) is modelled using a gamma distribution with parameters α = 0.8 and β = 0.4. . (c) Consider the overall rainfall in 365 days, and use mgfs and their properties to prove that this is Ga (292, 0.4). .
Answer #1

Let Xi denotes the rainfall in Dubin (measured in millimeters) of i-th day.

Here,

X_i \sim Gamma(\alpha = 0.8, \beta = 0.4)

The probability density function of X is

f(x) = \frac{0.4^0.8}{0.8!}e^{-0.4x}x^{0.8 - 1},x>0

The MGF of X is

E(e^{tX}) = \int_{0}^{\infty }e^{tx}f(x) dx= \int_{0}^{\infty }e^{tx}*\frac{0.4^{0.8}}{\Gamma (0.8)}e^{-0.4x}x^{0.8 - 1}dx

\Rightarrow E(e^{tX}) = \frac{0.4^{0.8}}{\Gamma (0.8)}\int_{0}^{\infty }e^{-(0.4-t)x}x^{0.8 - 1}dx

\Rightarrow E(e^{tX}) = \frac{0.4^{0.8}}{\Gamma (0.8)}\int_{0}^{\infty }e^{-(0.4-t)x}x^{0.8 - 1}dx

\Rightarrow E(e^{tX}) = \frac{0.4^{0.8}}{\Gamma (0.8)}*\frac{\Gamma (0.8)}{(0.4 -t)^{0.8}}

\Rightarrow E(e^{tX}) = \frac{0.4^{0.8}}{(0.4 -t)^{0.8}}

\Rightarrow E(e^{tX}) = \frac{1}{(1 -\frac{t}{0.4})^{0.8}}

Let us denote S = \sum_{i = 1}^{365}X_i

The MGF of S will be

M_S(t) = E(e^{tS}) = E(e^{t\sum_{i = 1}^{365}X_i}) =E(e^{tX_1})*E(e^{tX_2})*....*E(e^{tX_{365}}) since Xi 's are independent.

\Rightarrow M_S(t) =(\frac{1}{(1 -\frac{t}{0.4})^{0.8}})*(\frac{1}{(1 -\frac{t}{0.4})^{0.8}})*....*(\frac{1}{(1 -\frac{t}{0.4})^{0.8}})

\Rightarrow M_S(t) =(\frac{1}{(1 -\frac{t}{0.4})^{0.8}})^{365}

\Rightarrow M_S(t) =\frac{1}{(1 -\frac{t}{0.4})^{0.8*365}}

\Rightarrow M_S(t) =\frac{1}{(1 -\frac{t}{0.4})^{292}}

Since MGF uniquely identifies a distribution, so we can say

S \sim Gamma(\alpha = 292, \beta = 0.4)